OpenStudy (anonymous):

find all tangent lines to y = cubed root of x that have a slope of 3/4

OpenStudy (freckles):

so you know you need to solve the equation y'=3/4 for x?

OpenStudy (freckles):

derivative you should think slope of tangent line that is why you are solving y'=3/4

OpenStudy (anonymous):

I'm not getting it.

OpenStudy (freckles):

you are given the slope of the tangent line is 3/4 you know to find the slope of the tangent line you must find the derivative?

OpenStudy (anonymous):

Yes

OpenStudy (freckles):

so what is the problem?

OpenStudy (freckles):

differentiating y ? or what is it that you aren't getting in?

OpenStudy (freckles):

you can write your y as x^(1/3) and use power rule

OpenStudy (freckles):

\[(x^n)'=n x^{n-1}\]

OpenStudy (freckles):

\[(x^\frac{1}{3})'=?\]

OpenStudy (anonymous):

(x/3)^-2/3

OpenStudy (freckles):

the 1/3 part shouldn't be raised to the (-2/3)

OpenStudy (anonymous):

Yeah, I didn't intend that.

OpenStudy (freckles):

\[(x^n)'=nx^{n-1} \\ (x^\frac{1}{3})'=\frac{1}{3}x^{\frac{1}{3}-1}=\frac{1}{3}x^{\frac{-2}{3}}\]

OpenStudy (freckles):

\[y=x^\frac{1}{3} \\ y'=\frac{1}{3 x^{\frac{2}{3}}} \\ \text{ so remember we needed to solve } y'=\frac{3}{4} \text{ for } x \\ \text{ so we need to solve } \frac{1}{3x^\frac{2}{3}}=\frac{3}{4} \text{ for } x \]

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

i don't understand the y'=3/4

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

It just clicked

OpenStudy (freckles):

to find for which x values on the graph of f(x)=x^(1/3) we have the slopes of their tangent lines is 3/4

OpenStudy (anonymous):

I understand now

OpenStudy (freckles):

to find the tangent line we just need the slope and the point we already have the slope we just the need the point and remember the tangent line will look like this: \[y-f(a)=f'(a)(x-a) \\ \text{ where we already know } f'(a) \text{ to be } \frac{3}{4} \\ y-f(a)=\frac{3}{4}(x-a)\]

OpenStudy (freckles):

so that is what your answer will look like and you should have two answers

OpenStudy (anonymous):

Ok, how do I find X

OpenStudy (freckles):

you still having trouble solving that equation for x one sec let me repost it again

OpenStudy (freckles):

\[y=x^\frac{1}{3} \\ y'=\frac{1}{3 x^{\frac{2}{3}}} \\ \text{ so remember we needed \to solve } y'=\frac{3}{4} \text{ for } x \\ \text{ so we need \to solve } \frac{1}{3x^\frac{2}{3}}=\frac{3}{4} \text{ for } x \\ 3x^\frac{2}{3}=\frac{4}{3} \text{ just flipped both sides } \\ x^\frac{2}{3}=\frac{4}{9} \text{ just divided both sides by } 3 \\ \text{ you have even power so you will have two answers here } \\ x= \pm (\frac{4}{9})^\frac{3}{2}\] this answer can be simplified a bit

OpenStudy (freckles):

I suggest doing the square root part before doing the cubing

OpenStudy (anonymous):

So what is y1

OpenStudy (freckles):

y1? are you talking about y' ?

OpenStudy (freckles):

or are you talking about the corresponding y values for the x values for which the tangent lines with slope 3/4 occurs?

OpenStudy (anonymous):

In point slop formula the y1 as in y - y1 = m(x-x1)

OpenStudy (freckles):

ok so have are having trouble pluggin into y=x^(1/3) ?

OpenStudy (freckles):

you should have gotten the x values are -8/27 and 8/27

OpenStudy (freckles):

\[y=(\pm \frac{8}{27})^\frac{1}{3}=\pm (\frac{8}{27})^\frac{1}{3}\] just take the cube root of the numerator and denominator

OpenStudy (freckles):

\[y=\pm \frac{2}{3}\]

OpenStudy (freckles):

so tangent lines with slope 3/4 will go through the following two points on y=f(x): (-8/27,-2/3) (8/27,2/3)

OpenStudy (freckles):

and I don't mean the same tangent line will go through those two points I'm talking about two separate points on the curve which means we are talking about two separate tangent lines that satisfies the given conditions

OpenStudy (anonymous):

ok so y - 2/3 = 3/4(x- (-8/27)) y + 2/3 = 3/4(x+ (-8/27))

OpenStudy (freckles):

well m=3/4 with as you called it (x1,y1)=(-8/27,-2/3) means you have y+2/3=3/4(x+8/27) so this line is correct but m=3/4 with (x1,y1)=(8/27,2/3) gives y-2/3=3/4(x-8/27)

OpenStudy (freckles):

it was probably just a type-o you made and that is what you meant though

OpenStudy (anonymous):

Yep. Thanks again!

OpenStudy (freckles):

i didn't actually see the type-o in your second equation until just now... For some reason I only noticed the first type-o in the first equation.

OpenStudy (freckles):

anyways goodnight